OpenStudy (zenmo):

Find the given derivative by finding the first few derivatives and observing the pattern that occurs.

OpenStudy (zenmo):

\[\frac{ d ^{35} }{ dx ^{35} }(xsinx)\]

OpenStudy (zenmo):

Another example from the book: \[\frac{ d ^{99} }{ dx ^{99} }\] \[\frac{ d ^{1} }{ dx ^{1} }(sinx)=cosx\] \[\frac{ d ^{2} }{ dx ^{2} }(sinx)=-sinx\] \[\frac{ d ^{3} }{ dx ^{3} }(sinx)=-cosx\] \[\frac{ d ^{4} }{ dx ^{4} }(sinx)=sinx\] The derivatives of sinx occur in a cycle of four. Since 99 = 4(24) + 3, we have \[\frac{ d ^{99} }{ dx ^{99} }(sinx) = \frac{ d ^{3} }{ d x^{3} }(sinx) = -cosx\]

zepdrix (zepdrix):

Ya, your problem will work out similarly I think. But you have product rule, so uhhhh try a few, see what it looks like :)

OpenStudy (zenmo):

\[y' = (x)(cosx) + (1)(sinx) = xcosx + sinx\]

zepdrix (zepdrix):

So you chain ruled, what happened to the sinx on the far right though? Hmm

OpenStudy (zenmo):

are the derivatives of the first two correct?

zepdrix (zepdrix):

first derivative is good.

OpenStudy (zenmo):

\[y'' = (x)(-sinx) + (1)(cosx) + cosx = -xsinx + 2cosx\]

OpenStudy (zenmo):

Is the 2nd deriv. correct?

zepdrix (zepdrix):

Ok that looks better :)

zepdrix (zepdrix):

xsinx is what we started with, right?

OpenStudy (zenmo):

yes

zepdrix (zepdrix):

\(\large\rm y''=-y+2cos x\) I'm not sure if this helps us or not, just something to think about.

zepdrix (zepdrix):

I was trying something else sneaky here :) \(\large\rm y'''=-y'-2sin x\) \(\large\rm y^{(4)}=-y''-2cos x=-(-y+2cos x)-2cos x\)

zepdrix (zepdrix):

So every `two derivatives`, we gain 2 more cosines, and the sign changes on everything.

zepdrix (zepdrix):

If we don't want to deal with the sign changes, we can count by 4's instead maybe. So um, \(\large\rm y^{(4)}=y-4cos x\) Then, \(\large\rm y^{(32)}=y-32cos x\) somethingggggg like that maybe get us on the right track?

OpenStudy (zenmo):

so, the 32th derivative would get us back to xsinx?

zepdrix (zepdrix):

\(\large\rm y^{(32)}=x\sin x-32\cos x\) Assuming I didn't make a boo boo somewhere, ya it seems that way.

OpenStudy (zenmo):

the -32cosx shouldn't be there, I think

OpenStudy (zenmo):

should be just only xsinx

zepdrix (zepdrix):

Were the 2cosx cancelling out each time? Maybe I messed up a sign earlier, lemme check. Cause I thought we were GAINING 2cosx each two derivatives.

OpenStudy (zenmo):

It could be some trig. identity somewhere in the middle

OpenStudy (zenmo):

but then, just an assumption

zepdrix (zepdrix):

\(\large\rm y=x sin x\) \(\large\rm y'=x cos x+sin x\) \(\large\rm y''=-x sin x+2cos x=-y+2cos x\) \(\large\rm y'''=-y'-2sin x\) \(\rm y^{(4)}=-y''-2cos x=-(-x sin x+2cos x)-2cos x=x sin x-4cos x=y-4cos x\)

OpenStudy (zenmo):

Oh, you are actually correct

zepdrix (zepdrix):

y'' = -(y-2cosx) y^(4) = +(y-4cosx) Do you see how the sign is changing every two derivatives, and our cosines are increasing at the same rate as the derivative number?

OpenStudy (zenmo):

Seems like, every even derivative there is a cosine at the end. For every odd derivative there is a sine at the end

zepdrix (zepdrix):

I'm not sure if we should even worry about the odd derivatives :) If we look for a 5th derivative, maybe we can count by 5's.... But I'm thinking, using what we have here, we can get from the 4th to the 32nd very easily, and then from the 32nd to the 34th using what we know about the change that happens every 2 derivatives. And then take 1 derivative from there.

OpenStudy (zenmo):

\[\frac{ d }{ dx }=xcosx+sinx\] \[\frac{ d^{2} }{ dx^{2} }=-xsinx+2cosx\] \[\frac{ d^{3} }{ dx^{3} }=-xcosx-3sinx\] \[\frac{ d^{4} }{ dx^{4} }=xsinx-4cosx\] \[\frac{ d^{5} }{ dx^{5} }=xcosx+5sinx\] \[\frac{ d^{6} }{ dx^{6} }=-xsinx+6cosx\]

OpenStudy (zenmo):

The first 4 terms seem to be of the following choices: xcosx, -xsinx, -xcosx, xsinx and the last 4 terms: sinx, cosx, -sinx, -cosx

OpenStudy (zenmo):

\[\frac{ d^{35} }{ dx^{35} }=-xcosx-35sinx \] is the answer

zepdrix (zepdrix):

Is the xcosx positive maybe? Hmm sec I gotta check the pattern again :) lol

OpenStudy (zenmo):

yea, we got the math down, it is just analyzing the pattern

zepdrix (zepdrix):

nono I think you're right :) I was approaching it a little different though: y4 = y-4cosx So then y32 = y-32cosx and 2 more derivatives, y34 = -(y-34cosx) and doing the hard work taking another derivative, y35 = -(y'+34sinx) and then a little more work. Yah your way seems better, just looking at the pattern completely :D hehe

OpenStudy (zenmo):

One second.

OpenStudy (zenmo):

\[\frac{ d^{35} }{ dx^{35} }(xsinx) => 35 = 4(8) + 3 = 35\]

OpenStudy (zenmo):

Since, this one has 4 patterns (the same as the example)

zepdrix (zepdrix):

ooo neato :D

OpenStudy (zenmo):

so, it will look the same as \[\frac{ d^{3} }{ dx^{3} }\]

OpenStudy (zenmo):

-xcosx - sinx

zepdrix (zepdrix):

yah, very cool c:

OpenStudy (zenmo):

then, you just plug in the numbers

OpenStudy (zenmo):

-xcosx-35sinx

OpenStudy (zenmo):

:)

OpenStudy (zenmo):

That small "formula" helps, so we don't have to manually count